OpenStudy (anonymous):

An object is launched upward at 64 ft/sec from a platform that is 80 feet high. What is the objects maximum height if the equation of height (h) in terms of time (t) of the object is given by h(t) = -16t² + 64t + 80? Round to the nearest foot.

OpenStudy (anonymous):

So the maximum height of an object will be atteined at the time when the velocity of the object goes to zero. The velocity is the derivative of positions so,\[v(t)=(-16t^2+64t+80) \prime = -32t+64=0\] Solving for t, we see this happens at t=2seconds. Therefore, the height of the object at t=2 is:\[h(2)=-16*2^2+64*2+80=144feet\]

OpenStudy (anonymous):

so the answer is 144 feet